Web Assign and the result in your paper submission
Answer #1

R3 and R4 are in series and their combination is given as

R34 = R3 + R4 = 100 + 25 = 125 ohm

R34 and R2 are in parallel and their combination is given as

R234 = R2 R34 /(R2 + R34) = 25 x 125 /(25 + 125) = 20.83 ohm

R1 R234 and R5 are in series and their combination is given as

R12345 = R1 + R234 + R5 = 25 + 20.83 + 25 = 70.83 ohm

current in R1 and R5 = 6/70.83 = 0.085 A

V1 = V5 = 0.085 x 25 = 2.125

V2 = V34 = 6 - 2.125 - 2.125 = 1.75 volts

i3 = V34/R34 = 1.75/125 = 0.014 A